OpenStudy (anonymous):

a female lion had four cubs. they weighed 3 7/20, 3 9/20, 2 7/10, 2 4/5 whats the average cub weight?

OpenStudy (amistre64):

add em up ... 3 7/20, 3 9/20, 2 7/10, 2 4/5 = N 20(3 7/20, 3 9/20, 2 7/10, 2 4/5 = N) 60 +7+60+9+40+14+40+16 = 20N 180+18+80+14+16 = 20N 180+98+30 = 20N 210+98 = 20N 308 = 20N 154 = 10N 15.4 = N

OpenStudy (amistre64):

15.4 / 4 = 3.8

OpenStudy (amistre64):

but my head math migh tbe off

OpenStudy (anonymous):

3 7/20 + 3 9/20 + 2 14/20 + 2 16/20 = 12 6/20 = 12 3/10 (12 3/10) / 4 cubs = 3 3/40 on average.

OpenStudy (anonymous):

What amistre says is right, add all the values and then divide by number of cubs. However I think he made a summation mistake.

OpenStudy (amistre64):

yeah, the add up was spose to be 246, not 308 :)

OpenStudy (amistre64):

to wit, i would have gotten 12.3 .. and came to your results as well ;)

OpenStudy (amistre64):

60+60 = 120 , not 180 .... i forgot how to add lol